dinonuggets12:

The commutative property of addition says you can add numbers in any order and get the same answer. Is the same true for subtraction? 22 – 12 10 12 – 22 –10 10 ≠ –10 22 – 12 ≠ 12 – 22 a – b ≠ b – a Changing the order in which you subtract numbers does change the result, so the commutative property does not work for subtraction.

dinonuggets12:

A. 2 – 12 = 12 – 2 B. (3 – 11) – 5 = 3 – (11 – 5) C. 8 + 7 = 7 + 8 D. (6 + 9) + 4 = 6 + (9 + 4)